a)x^2-2x=15
B)3x^2+19+1=0
thank you

Answer by stanbon(75887)

a)x^2-2x=15
x^2-2x-15=0
(x-5)(x+2)=0
x=5 or x=-2
-----------------
B)3x^2+19+1=0
Did you post what you wanted?
Cheers,
Stan H.